Kevin operates a soybean farm. He buys many supplies in bulk. Often the bulk products need to be custom mixed before Kevin can use them. To apply herbicide to a large field he must mix a solution of 67% herbicide with a solution of 46% herbicide to form 42 liters of a 55% solution. How much of the 67% solution must he use? Write a system of equations to find how much of the 67% solution must he use. Identify the variables, write the equations, and solve the system using Elimination or Substitution methods. Write complete Answers. Show all your work step by step. Write a short explanation about why you chose the best method to solve the problem.
SHOW ALLLL WORK PLEASE!!!!!!!!!!!!

Answer by checkley79(3341)

.67X+.46(42-X)=.55*42
.67X+19.32-.46X=23.1
.21X=23.1-19.32
.21X=3.78
X=3.78/.21
X=18 LITERS OF THE 67% MIX IS USED.
42-18=24 LITERS OF THE 46% MIX IS USED.
PROOF:
.67*18+.46*24=23.1
12.06+11.04=23.1
23.1=23.1
THIS IS THE STANDARD METHOD BECAUSE IT IS THE SHORTEST.